OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8 -HELP PLEASEEEEEEEEEEEEEEEEEEEEEEEEEE

OpenStudy (anonymous):

PLEASE HELP ME... xox

OpenStudy (anonymous):

anyone wanna help? xx

OpenStudy (anonymous):

pwease??????????????????????

OpenStudy (anonymous):

babes? help me

OpenStudy (anonymous):

you there! help me out

OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8 --------------------------------------------------------------------------------

OpenStudy (anonymous):

help me. I only have a few minutes to finish a test!

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]--> use the equ

OpenStudy (anonymous):

equation of the line..

OpenStudy (anonymous):

slope of parallel line has same slope with the original equation.. m=slope y=mx+b-->original equation

OpenStudy (anonymous):

i don't understand... can you explain it to meeeeeeeeeeeee

OpenStudy (anonymous):

first you need to find the slope, slope of parallel line has same slope with the original equation.. so m(slope) is ___?

OpenStudy (anonymous):

(6,-2) ?

OpenStudy (anonymous):

y = x + 7. -------> original function in the for y=mx+b, where m is the slope and b is the y-int. now tell me, what will be the slope?

OpenStudy (anonymous):

I have no clue..... 7? or do you mean whats the slope from the original equation?

OpenStudy (anonymous):

yes, the slope of the given equation.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):